OpenStudy (anonymous):

how would i find the slope that is perpendicular to a line that passes through (2,4) and (8,1)

OpenStudy (zzr0ck3r):

what is the slope of the line that goes through (2,4) and (8,1)?

OpenStudy (zzr0ck3r):

if we have two points \[(x_0,y_0) \space and \space (x_1,y_1)\] then the slope is \[slope = \frac{y_1-y_0}{x_1-x_0}\]

OpenStudy (zzr0ck3r):

so what is the slope @istinkatmath1 ?

OpenStudy (anonymous):

the slope would equal to two

OpenStudy (zzr0ck3r):

(2,4) and (8,1) \[\frac{1-4}{8-2}=?\]

OpenStudy (anonymous):

3/-6 then you simplify and you get 2

OpenStudy (anonymous):

Two lines are perpendicular if the product of their slopes = -1

OpenStudy (anonymous):

So if you know the slope of a given line, the slope of the line perpendicular to it is the negative reciprocal

OpenStudy (anonymous):

-3/6 = - 1/2 So it's negative reciprocal is 2

OpenStudy (anonymous):

okay thank you so much!

OpenStudy (zzr0ck3r):

\[\frac{-3}{6}=\frac{-1}{2}\ne2\]

OpenStudy (anonymous):

Who is saying they are equal? -1/2 and 2 are negative reciprocals. In other words when you multiply them the product is -1. That will now mean the lines are perpendicular